OpenStudy (maddiie):

Find the soultion to each system by the addition(elimination) method. 4s-5t=18 6s + 10t=-8

OpenStudy (anonymous):

multiply first equation by 2\[8s-10t=36\] \[6s+10t=-8\]

OpenStudy (anonymous):

add to get \[14s=28\] \[s=2\]

OpenStudy (maddiie):

5(6s)+5(10t)=5(-8) 30s+50t=-40

OpenStudy (anonymous):

(4s-5t=18 )*2 ---->8s-10t=36 6s + 10t=-8, then add 14s=28 s=2 4(2)-5t=18 8-5t=18 -5t=18-8 -5t=10 t=-2

OpenStudy (anonymous):

now substitute back to find t

OpenStudy (anonymous):

(4s-5t=18)2 6s+10t=-8 ____________ 8s-10t=36 6s+10t=-8 ____________ 14s = 28 s= 2 4(2)-5t=18 8-18=5t -10=5t t= -2

OpenStudy (anonymous):

$$ \int \sum \text{x} \partial{} $$

OpenStudy (anonymous):

i'mjk

OpenStudy (anonymous):

you ecin. hmm you love calculus?

OpenStudy (anonymous):

ok what does this mean The ratio of the radii of two circles at the centers of which two arcs of the same length subtended a angles of 60 degree and 70 degree is

OpenStudy (anonymous):

i wrote sex ...

OpenStudy (amistre64):

blasphemy! ..... maybe lol

OpenStudy (anonymous):

sex? that is bad

OpenStudy (anonymous):

we can use a formula there satellite

OpenStudy (maddiie):

Guys go chat on the group chat i'm trying to get help.

OpenStudy (anonymous):

(4s-5t=18 )*2 ---->8s-10t=36 6s + 10t=-8, then add 14s=28 s=2 4(2)-5t=18 8-5t=18 -5t=18-8 -5t=10 t=-2 posted by pat18

OpenStudy (anonymous):

ecin?

OpenStudy (maddiie):

Ok i understood. I hate math

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

i love math